Ocean waves are traveling to the east at 3.9 m/s with a distance of 21 m between crests. (a) With what frequency do the waves hit the front of a boat when the boat is at anchor? Hz (b) With what frequency do the waves hit the front of a boat when the boat is moving westward at 1.5 m/s? Hz
Answer #1


v = 3.9m/s ? = 21

a) v = ?*f


f = 3.9/21 = 0.186 Hz

b) f' = (3.9+1.5)/21 = 0.257 Hz
Answer #2

v = 3.8m/s     ?x = 17m = (?)

a. v = F?

3.8m/s = F (1.7m)

F = (3.8m/s) / (17m) = 0.22 Hz Anchored

b. v = 1.0m/s   ?x = 17m = ( ?) Use same formula as above.

F = (1.0m/s) / (17m) = 0.059Hz moving westward

Add the two together for answer 0.22Hz + 0.059 = 0.279Hz